# Oracle-based Uniform Sampling from Convex Bodies

### v1

@@ Abstract
We propose new Markov chain Monte Carlo algorithms to sample a uniform distribution on a convex body $K$. Our algorithms are based on the Alternating Sampling Framework/proximal sampler, which uses Gibbs sampling on an augmented distribution and assumes access to the so-called restricted Gaussian oracle (RGO). The key contribution of this work is the efficient implementation of RGO for uniform sampling on $K$ via rejection sampling and access to either a projection oracle or a separation oracle on $K$. In both oracle cases, we establish non-asymptotic complexities to obtain unbiased samples where the accuracy is measured in Rényi divergence or $\chi^{2}$-divergence.
Key words. Uniform sampling, Markov chain Monte Carlo, Alternating Sampling Framework, restricted Gaussian oracle, projection oracle, separation oracle, rejection sampling.
@@ 1 Introduction
Sampling points from convex bodies in high dimension is a classical and central problem in computational geometry, probability, statistics, and optimization. Given a convex body $K\subseteq\mathbb{R}^{d}$, one likes to generate samples according to some distributions defined on $K$. Past and recent works in the area of constrained sampling in high dimension include [52, 12, 30, 4, 3, 31, 36, 21, 44, 35, 19, 66, 28, 1, 47, 58, 46, 24, 23]; and many others. In this paper, we will focus on uniform sampling on $K$ which can be viewed as the most fundamental case. It is closely related to the problem of efficiently computing the volume of $K$, which is a important problem in computer science for the last few decades (see [13] and the references therein). Moreover, uniform sampling also has a connection to Bayesian inference. If one takes the Gaussian distribution $\mathcal{N}(0,\sigma^{2}I_{d})$, restricts it to $K$ and lets $\sigma$ becomes sufficiently large, then this truncated Gaussian distribution resembles the uniform distribution on $K$. At the same time, truncated Gaussian distribution has been used extensively in Bayesian statistical models with probit regression and censored data, see [25, 2, 10, 26, 59]. Due to its importance, many works have been devoted to develop algorithms for this problem. The seminal work by [16] proposes the first algorithm to approximate the volume of any convex $K$ in polynomial time and also introduces the celebrated Ball walk to generate samples uniformly on $K$. Assuming a membership oracle, an iteration of the Ball walk works as follows:
  • pick a uniform random point $y$ from the ball of radius $\delta$ centered at the current point $x$;
  • if $y$ is in $K$, go to $y$; otherwise stay at the current point $x$.
Other algorithms that can be used for uniform sampling from convex $K$ are the Hit-and-Run walk in [57, 53, 51], the coordinate Hit-and-Run walk [60, 14, 15], the Dikin walk in [31], the Geodesic walk in [43], and diffusion-based samplers in [4, 3, 24, 5], among others. More details on these algorithms are provided in Appendix C.
@@
Related works. In [42], Lee, Shen, and Tian develop a Gibbs sampling scheme for log-concave sampling in high dimension. They name it Alternating Sampling Framework (ASF), which is also referred to as the proximal sampler due to its connection to the proximal point method in optimization. Each iteration of the scheme consists of an initial Gaussian step followed by a proximal-type sampling step. The ASF is interesting to study in several ways. First, it is at the intersection of sampling and optimization. Second, it is a high-accuracy/non-biased sampler (compared to Langevin Monte Carlo/Underdamped Langevin Monte Carlo, which are known to be biased). Finally, thanks to the discovery in [8], the ASF as a discrete-time Markov chain can be viewed through the lens of Itô diffusion processes and analyzed using tools from stochastic calculus. For these reasons, the ASF/proximal sampler inspires many follow-up works such as [22, 7, 54, 63, 65, 64, 48, 18, 49] among others.
In [39], Kook, Vempala, and Zhang propose the In-and-Out algorithm to perform uniform sampling from convex $K$. Their algorithm is based on the ASF and they employ a clever smoothing argument to adapt the proof technique in [8] to the case of the uniform distribution from $K$. In particular, their implementation of the RGO step is via a membership oracle: they sample $x$ from a Gaussian distribution until $x\in K$ up to a certain number of maximum attempts $N$, at which point the algorithm halts and declares failure. Then they carefully analyze the condition on step size and $N$ to make sure the failure probability is small. Subsequent works using ASF for uniform sampling on $K$ and more general log-concave sampling by reducing it to the problem of exponential sampling from convex bodies have been carried out in [38, 37, 34]. In particular, [34] is able to get rid of the failure probability in the In-and-Out algorithm by introducing a restart step: if one cannot generate $x$ that is in $K$ from a Gaussian distribution after $N$ attempts in the second step of the ASF, then restart by returning to the first step of the ASF.
@@
Our contributions. We develop efficient algorithms to perform uniform sampling on the convex body $K$. Our algorithms are also based on the ASF/proximal sampler like the In-and-out algorithm in [39], however, our goal is to explore beyond the membership oracle employed in [39] and to consider other common oracles in convex optimization and computer science, which are projection oracles and separation oracles. More specifically, each iteration of the proximal sampler consists of an initial Gaussian step followed by a proximal-type sampling step. The latter is known as the Restricted Gaussian Oracle (RGO) and is the primary challenge of applying the ASF. We propose Algorithm 3 and Algorithm 4 as implementations of the RGO via rejection sampling. They respectively require a projection oracle and a separation oracle on $K$. An advantage of our algorithms is that our RGO implementations are unbiased and the outputs of our RGO implementations belong to the feasible set $K$ almost surely (see Remark 3.1). Therefore, we are able to avoid the failure probability encountered in [39, 38, 37] and thus offer another alternative to the restart procedure in [34].
@@
More specifically, in Section 3, we assume the convex set $K$ satisfies $B(0,1)\subseteq K\subseteq B(0,R)$ and the initial distribution $\mu_{0}$ of the ASF (Algorithm 2) satisfies a warm-start assumption: $d\mu_{0}/d\pi\leq M$ where $\pi\sim\mathbf{1}_{K}$. Moreover, the RGO implementation is via Algorithm 3 which uses a projection oracle on $K$ and rejection sampling. Then per our Theorem 3.5, to achieve an $\epsilon$-accuracy in Rényi divergence $\mathcal{R}_{q}$, Algorithm 2 combined with Algorithm 3 needs at most
$\mathcal{O}\left(d^{2}{C_{\mathrm{LSI}}q\log\left(2\frac{\log M}{\epsilon}\right)}\right).$ (1)
@@
iterations. $C_{\mathrm{LSI}}$ as the LSI constant of the uniform distribution on $K$ is of the order $\mathcal{O}(D^{2})$ where $D$ is the diameter of $K$. Our result therefore matches the iteration complexity (in term of dimension dependence and step size dependence) of the In-and-out algorithm [39, Theorem 27] and that of the Ball walk (see our Appendix C). Moreover, under the additional assumption that the step size $\eta$ of the ASF equals $1/d^{2}$, Theorem 3.5 also says that each iteration of the ASF makes one query to the projection oracle on $K$ and has at most an average of $M(\sqrt{2\pi e}+1)$ rejections for the rejection sampling. Finally, Theorem 3.5 also offers similar results in $\chi^{2}$-divergence.
@@
In Section 4, in the absence of a projection oracle on $K$, we implement the RGO via Algorithm 4 which uses a separation oracle on $K$ and rejection sampling. We assume the same conditions about $K$ as in the previous paragraph. Then Theorem 4.3 says that the number of iterations of the ASF (Algorithm 2) to reach $\epsilon$-accuracy in Rényi divergence $\mathcal{R}_{q}$ does not exceed the value at (1). In particular, each iteration of the ASF makes ${\cal O}\left(d\log\frac{d\gamma}{\alpha}\right)$ queries to the separation oracle on $K$, where $\gamma=R/\mbox{minwidth}(K)$, $\mbox{minwidth}(K)=\min_{\left\lVert a\right\rVert=1}\left(\max_{y\in K}a^{T}y-\min_{y\in K}a^{T}y\right)$, and the constant $\alpha\in(0,1)$ is not too small (in the sense that it satisfies $\Pr\left(\alpha\leq\frac{2}{d^{3}7R^{2}}\right)\leq 4\exp\left(-\frac{d^{2}R^{2}}{8}\right)$). In addition, the average number of rejections is $M\sqrt{2\pi}\exp\left(\frac{13}{4}+\frac{20}{d}\right)+M\exp\left(\frac{9}{4}+\frac{12}{d}\right)$ for the rejection sampling. Finally, Theorem 4.3 also provides similar results in $\chi^{2}$-divergence.
@@ 2.1 Notation, definitions, and assumptions
Regarding notation, $\left\lVert\cdot\right\rVert$ denotes the Euclidean norm on $\mathbb{R}^{d}$, $\left\lVert\cdot\right\rVert_{\operatorname{op}}$ denotes the matrix operator norm, and $I_{d}$ denotes the identity matrix of size $d\times d$. The expression $x=\mathcal{O}(a)$ means there exists a universal constant $C>0$ such that $x\leq Ca$. The notation $\tilde{\cal O}(a)$ is used similarly, but allows additional logarithmic factors.
Absolute continuity. Assume $\mu,\nu$ are two measures on a probability space $(E,\mathcal{F})$. We say $\mu$ is absolutely continuous with respect to $\nu$, denoted by $\mu\ll\nu$, if there exists a function $f:E\to\mathbb{R}$ such that for any $A\in\mathcal{F}$, we have $\mu(A)=\int_{A}f(x)d\nu(x)$. The function $f$ is called the Radon-Nikodym derivative of $\mu$ with respect to $\nu$, and is denoted as $\frac{d\mu}{d\nu}$.
Metric. Assume that $\phi$ is a convex function $\mathbb{R}_{\geq 0}\to\mathbb{R}$ with $\phi(1)=0$ and that $\mu,\nu$ are two probability measures on a probability space $(E,\mathcal{F})$ satisfying $\mu\ll\nu$, then the $\phi$-divergence between $\mu$ and $\nu$ on $\mathbb{R}^{d}$ is defined as
$D_{\phi}(\mu||\nu)=\int_{E}\phi\left(\frac{d\mu}{d\nu}\right)d\nu.$
When $\phi(x)=x\log x$, this is the Kullback-Leiber divergence and when $\phi(x)=x^{2}-1$, this is the $\chi^{2}$-divergence. Moreover, for $q>0$, the $q$-Rényi divergence is
$\mathcal{R}_{q}(\mu||\nu)=\frac{1}{q-1}\log\left(\chi^{q}(\mu||\nu)+1\right).$
Finally, the relative Fisher information is defined as
$\mathrm{FI}(\mu||\nu)=\int_{E}\left\lVert\nabla\log\frac{d\mu}{d\nu}\right\rVert^{2}d\mu.$
Isoperimetric inequalities and isoperimetric constants. Regarding isoperimetric inequalities, we say $\nu$ satisfies the log-Sobolev inequality (LSI) with constant $C_{\mathrm{LSI}}$ if for all $\mu\ll\nu$,
$\mathrm{KL}(\mu||\nu)\leq\frac{C_{\mathrm{LSI}}}{2}\mathrm{FI}(\mu||\nu).$
Meanwhile, we say $\nu$ satisfies the Poincaré inequality (PI) with constant $C_{\mathrm{PI}}$ if for any smooth bounded function $\psi$,
$\mathsf{Var}_{\nu}\left(\psi\right)\leq C_{\mathrm{PI}}\mathbb{E}_{\nu}\left[\left\lVert\nabla\psi\right\rVert^{2}\right].$
In [39, Appendix C], the authors provide a nice summary of studies on isoperimetric constants in [6, 29, 45, 9, 33, 32]. A consequence of the aforementioned studies is the following result about LSI and PI constants of the uniform distribution on convex $K$. We note that a probability measure $\pi$ on $K$ is isotropic if for a random vector $(X_{1},\ldots,X_{d})$ distributed as $\pi$, we have $\mathbb{E}{X_{i}}=0$ and $\mathbb{E}{X_{i}X_{j}}=\mathbf{1}_{i=j}$ for all $1\leq i,j\leq d$.
@@
[39, Lemma 18] Let $\pi$ be the uniform distribution over $K$ and $K\subset\mathbb{R}^{d}$ be a convex body with diameter $D$, where $D=\max_{x,y\in K}\left\lVert x-y\right\rVert$. Then we have $C_{\mathrm{PI}}(\pi)=\mathcal{O}\left(\left\lVert\mathrm{Cov}{\left(\pi\right)}\right\rVert_{\operatorname{op}}\log d\right)$ and $C_{\mathrm{LSI}}(\pi)=\mathcal{O}\left(D^{2}\right)$.
In particular, if $\pi$ is isotropic, then $C_{\mathrm{PI}}(\pi)=\mathcal{O}\left(\log d\right)$ and $C_{\mathrm{LSI}}(\pi)=\mathcal{O}\left(D\right)$.
Oracles. For any query point $x\in\mathbb{R}^{d}$, a membership oracle on $K$ provides the correct answer to whether $x\in K$. Meanwhile, for any query point $x\in\mathbb{R}^{d}$, a separation oracle on $K$ either confirms $x\in K$ or if $x\notin K$, returns $g:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ satisfying for every $y\in K$,
$\langle g(x)\,,x-y\rangle\geq 0.$
One can immediately see that a separation oracle assumes a membership oracle. Finally, a projection oracle on $K$ provides the vector $\operatorname{proj}_{K}(y)=\mathrm{argmin}\,\{\left\lVert x-y\right\rVert^{2}:x\in K\}$ for any query point $y\in\mathbb{R}^{d}$. Clearly, $\operatorname{proj}_{K}(y)=y$ if $y\in K$.
Volumes. Let $\mathsf{vol}(K)$ and $\mathsf{vol}_{d-1}(\partial K)$ respectively denote the volumes of $K\subseteq\mathbb{R}^{d}$ and the boundary set $\partial K\subseteq\mathbb{R}^{d-1}$.
Warmness. Given probability measures $\mu,\nu$ on $\mathbb{R}^{d}$ and $M>0$, we says $\mu$ is $M$-warm with respect $\nu$ if $\mu\ll\nu$ and $\frac{d\mu}{d\nu}(x)\leq M,\forall x\in\mathbb{R}^{d}$. In both Algorithm 3 and Algorithm 4 of the paper, we will assume a warm start assumption: the starting distribution $\mu_{0}$ is $M$-warm with respect to the uniform distribution on $K$.
Standing assumptions for the paper. In both Sections 3 and 4, we assume the following hold:
@@
    the initial distribution $\mu_{0}$ is $M$-warm with respect to the uniform distribution on $K$, i.e., $\frac{d\mu_{0}}{d\pi}\leq M$ where $\pi\sim\mathbf{1}_{K}$ (warm-start assumption).
2.2 Alternating sampling framework
ASF is first proposed in [42] to sample log-concave distribution in $\mathbb{R}^{d}$. It is closely related to the proximal point method in optimization and is therefore also known as the proximal sampler. Given a step size $\eta>0$, it aims to sample the target distribution $\pi^{X}(x)\sim\exp(-f(x))$ by performing Gibbs sampling for the augmented distribution $\pi^{X,Y}(x,y)\sim\exp\left(-f(x)-\frac{\left\lVert x-y\right\rVert^{2}}{2\eta}\right)$ whose $X$-marginal is the target $\pi^{X}$. This idea of sampling from a joint distribution to obtain the marginal distribution has been observed in earlier references, for example [12]. Each ASF iteration alternates between two steps:
Algorithm 1 Alternating Sampling Framework [42]

### v2

@@ Abstract
We propose new Markov chain Monte Carlo algorithms to sample a uniform distribution on a convex body $K$. Our algorithms are based on the proximal sampler, which uses Gibbs sampling on an augmented distribution and assumes access to the so-called restricted Gaussian oracle (RGO). The key contribution of this work is an efficient implementation of the RGO for uniform sampling on convex $K$ that goes beyond the membership-oracle model used in many classical and modern uniform samplers, and instead leverages richer oracle access commonly assumed in convex optimization. We implement the RGO via rejection sampling and access to either a projection oracle or a separation oracle on $K$. In both oracle models, we provide non-asymptotic complexity guarantees for obtaining unbiased samples, with accuracy quantified in Rényi divergence and $\chi^{2}$-divergence, and we support these theoretical guarantees with numerical experiments.
Key words. Uniform sampling, Markov chain Monte Carlo, proximal sampler, restricted Gaussian oracle, projection oracle, separation oracle, rejection sampling.
@@ 1 Introduction
Sampling points from convex bodies in high dimension is a classical and central problem in computational geometry, probability, statistics, and optimization. Given a convex body $K\subseteq\mathbb{R}^{d}$, one likes to generate samples according to some distributions defined on $K$. Past and recent works in the area of constrained sampling in high dimension include [58, 16, 36, 6, 4, 37, 42, 27, 50, 40, 25, 77, 34, 1, 54, 68, 53, 30, 28]; and many others. In this paper, we will focus on the most fundamental case, i.e., uniform sampling on $K$. It is closely related to the problem of efficiently computing the volume of $K$, which is an important problem in computer science for the last few decades, see [17] and references therein. Moreover, uniform sampling also has a connection to Bayesian inference. If one takes the Gaussian distribution $\mathcal{N}(0,\sigma^{2}I_{d})$, restricts it to $K$ and lets $\sigma$ becomes sufficiently large, then this truncated Gaussian distribution resembles the uniform distribution on $K$. At the same time, truncated Gaussian distribution has been used extensively in Bayesian statistical models with probit regression and censored data, see [31, 2, 13, 32, 69]. Due to its importance, many works have been devoted to developing algorithms for this problem. The seminal work by [22] proposes the first algorithm to approximate the volume of any convex $K$ in polynomial time and also introduces the celebrated Ball walk to generate samples uniformly on $K$. Other algorithms that can be used for uniform sampling from convex $K$ are the Hit-and-Run walk in [67, 59, 61], the coordinate Hit-and-Run walk [71, 19, 20], the Dikin walk in [37], the Geodesic walk in [49], and diffusion-based samplers in [6, 4, 30, 8], among others. More details on these algorithms are provided in Appendix A.
@@
In [48], Lee, Shen, and Tian develop a Gibbs sampling scheme for log-concave sampling in high dimension. It is named Alternating Sampling Framework, which is commonly referred to as the proximal sampler. Each iteration of the scheme consists of an initial Gaussian step followed by a proximal-type sampling step. The proximal sampler is interesting to study in several ways. First, its Restricted Gaussian Oracle step is the sampling analogue of the proximal map in the proximal point method from optimization. Second, it is a high-accuracy/unbiased sampler (compared to Langevin Monte Carlo/Underdamped Langevin Monte Carlo, which are known to be biased). Finally, thanks to the discovery by [11], the proximal sampler as a discrete-time Markov chain can be viewed through the lens of Itô diffusion processes and analyzed using tools from stochastic calculus. For these reasons, the proximal sampler inspires many follow-up works such as [29, 10, 62, 74, 76, 75, 55, 24, 56] among others.
@@
Our contributions. We develop efficient algorithms to perform uniform sampling on the convex body $K$, and we complement our theoretical guarantees with numerical experiments. Our algorithms are based on the proximal sampler. A central goal of our work is to incorporate more ideas from modern optimization into the design and analysis of sampling algorithms. In particular, we move beyond the classical membership-oracle model, which underlies many random-walk-based samplers including the In-and-Out algorithm by [41], and instead explore richer oracle structures such as projection oracles and separation oracles, which are standard oracles in convex optimization. More specifically, each outer iteration of the proximal sampler consists of an initial Gaussian step followed by a proximal-type sampling step. The latter is known as the Restricted Gaussian Oracle (RGO) and is the primary challenge in applying the proximal sampler. We propose Algorithm 3 and Algorithm 4 as implementations of the RGO via rejection sampling; they respectively require a projection oracle and a separation oracle on $K$.
We summarize our main theoretical results as follows.
@@
  • Projection-oracle-based proximal sampler. In Section 3, we assume that $K$ satisfies $B(0,1)\subseteq K\subseteq B(0,R)$ where $B(0,R)$ is an $\ell_{2}$-ball centered at 0 with radius $R\geq 1$, and that the initial distribution $\mu_{0}$ of the proximal sampler (Algorithm 2) satisfies a warm-start assumption: $d\mu_{0}/d\pi\leq M$ where $\pi$ is the uniform distribution on $K$. We implement the RGO via Algorithm 3, which uses a projection oracle on $K$ and rejection sampling. Then, by Theorem 2.2 and Theorem 3.5, to achieve $\epsilon$-accuracy in Rényi divergence $\mathcal{R}_{q}$, Algorithm 2 combined with Algorithm 3 requires at most
    $\mathcal{O}\left(d^{2}{C_{\mathrm{LSI}}q\log\left(2\frac{\log M}{\epsilon}\right)}\right)$ (1)
@@
    proximal sampler iterations. Here $C_{\mathrm{LSI}}$, the LSI constant of the uniform distribution on $K$, is of order $\mathcal{O}(D^{2})$ where $D$ is the diameter of $K$. Our result therefore matches the iteration complexity (in terms of dimension dependence and stepsize dependence) of In-and-Out [41, Theorem 27] and that of the Ball walk (see Appendix A). Moreover, under the additional assumption that the stepsize $\eta$ of the proximal sampler equals $1/d^{2}$, Algorithm 3 and Theorem 3.5 also imply that each iteration of the proximal sampler makes one query to the projection oracle on $K$ and has at most an average of $M(\sqrt{2\pi e}+1)$ proposals for the rejection sampling. Finally, Theorem 2.2 and Theorem 3.5 also provide analogous results in $\chi^{2}$-divergence.
@@
  • Separation-oracle-based proximal sampler. In Section 4, we implement the RGO via Algorithm 4, which uses a separation oracle on $K$ and rejection sampling. We assume the same conditions on $K$ and $\mu_{0}$ as above. Then Theorem 2.2 and Theorem 4.2 imply that the number of iterations of the proximal sampler (Algorithm 2) to reach $\epsilon$-accuracy in Rényi divergence $\mathcal{R}_{q}$ has the same bound as in (1). In particular, each iteration of the proximal sampler makes ${\cal O}\!\left(d\log\frac{d\gamma}{\alpha}\right)$ queries to the separation oracle on $K$, where $\gamma=R/\mbox{minwidth}(K)$, $\mbox{minwidth}(K)=\min_{\left\lVert a\right\rVert=1}\left(\max_{y\in K}a^{T}y-\min_{y\in K}a^{T}y\right)$, and the constant $\alpha\in(0,1)$ is not too small, i.e., it satisfies $\Pr\left(\alpha\leq\frac{2}{7d^{3}R^{2}}\right)\leq 4\exp\left(-\frac{d^{2}R^{2}}{8}\right)$. In addition, the average number of proposals is $M\sqrt{2\pi}\exp\left(\frac{13}{4}+\frac{20}{d}\right)+M\exp\left(\frac{9}{4}+\frac{12}{d}\right)$ for the rejection sampling. Finally, Theorem 2.2 and Theorem 4.2 also provide analogous results in $\chi^{2}$-divergence.
Regarding the numerical experiments presented in Section 5, we compare our PS+Alg3 (Algorithm 2+Algorithm 3) with In-and-Out (Algorithm 1 in [41]) on random dense $Z$-polytopes, which are standard examples of convex bodies and are defined in Appendix D. We consider two stepsizes: the conservative choice $\eta_{A}=10^{-5}$ (as recommended by our Theorem 3.5) and an aggressive practical choice $\eta_{B}=10^{-2}$. As an accuracy proxy, we report the total variation (TV) error of the $x_{1}$ marginal against a long hit-and-run reference chain, and we also report elapsed time to capture solver/oracle overhead. Under the conservative stepsize, PS+Alg3 reduces TV faster and reaches a substantially lower final TV than In-and-Out, yielding a better accuracy-time tradeoff despite higher per-iteration cost. Under the aggressive stepsize, In-and-Out remains stable and achieves a final TV comparable to PS+Alg3, while PS+Alg3 reaches a given TV level substantially sooner in elapsed time.
Comparison to prior works.
To the best of our knowledge, there are two prior works on constrained sampling that design samplers using a projection oracle for the constraint set, namely [5, 52]. In particular, [52] studies sampling from a density proportional to $e^{-f}$ over a convex set $K\subset\mathbb{R}^{d}$, where $f$ is convex and Lipschitz continuous. They assume access to a projection oracle over $K$ and derive an iteration complexity bound of $O\!\left(\frac{d}{\epsilon^{2}}\right)$, where $d$ is the dimension and $\epsilon$ is the target accuracy in Wasserstein distance. Comparing the iteration complexity of [52] to (1), our guarantee is stated in a stronger metric (Rényi divergence and $\chi^{2}$ divergence), has better $\epsilon$-dependence but worse dependence on the dimension $d$.
Meanwhile, to the best of our knowledge, there is one earlier work by [36] that designs a constrained sampler under a separation-oracle model. The uniform-sampling algorithm therein has oracle complexity $O\!\left(d^{3}\log^{2}\!\frac{d}{\epsilon}\right)$ with respect to the total variation distance [36, Theorems 2.2 and 4.14]. In contrast, our proximal sampler with the RGO implementation given by Algorithm 4 achieves a comparable complexity $O\!\left(d^{3}\log d\cdot\log\frac{1}{\epsilon}\right)$ in stronger metrics, namely Rényi divergence and $\chi^{2}$ divergence.
It is also important to mention the recent work by [41]. Therein, Kook, Vempala, and Zhang propose the In-and-Out algorithm for uniform sampling from convex $K$. Their algorithm is based on the proximal sampler and they employ a clever smoothing argument to adapt the proof technique by [11] to the case of the uniform distribution from $K$. In particular, their implementation of the RGO step is via a membership oracle: they sample $x$ from a Gaussian distribution until $x\in K$ up to a certain number of maximum attempts $N$, at which point the algorithm halts and declares failure. Then they carefully analyze the condition on stepsize and $N$ to make sure the failure probability is small. Subsequent works using the proximal sampler for uniform sampling on $K$ and more general log-concave sampling by reducing it to the problem of exponential sampling from convex bodies have been carried out in [44, 43, 45]. In particular, [45] is able to get rid of the failure probability in the In-and-Out algorithm by introducing a restart step: if one cannot generate $x$ that is in $K$ from a Gaussian distribution after $N$ attempts in the second step of the proximal sampler, then restart by returning to the first step.
Finally, we highlight two additional features of our approach. First, our RGO implementations are unbiased and their outputs lie in $K$ almost surely (see Remark 3.1). Thus we avoid the failure probability in [41, 44, 43] and provide an alternative to the restart procedure of [45]. Second, to facilitate implementation, Appendix D details how to realize membership, projection, and separation oracles for common convex bodies. It turns out in many standard examples, these oracles have comparable computational cost to implement.
@@ 2.1 Notation, definitions, and assumptions
Regarding notation, $\left\lVert\cdot\right\rVert$ denotes the Euclidean norm on $\mathbb{R}^{d}$, $\left\lVert\cdot\right\rVert_{\operatorname{op}}$ the operator norm, and $I_{d}$ the $d\times d$ identity. The notation $x=\mathcal{O}(a)$ means $x\leq Ca$ for a universal constant $C>0$, and $\tilde{\mathcal{O}}(a)$ allows additional logarithmic factors.
Absolute continuity. For measures $\mu,\nu$ on $(E,\mathcal{F})$, we write $\mu\ll\nu$ if there exists $f:E\to\mathbb{R}$ such that $\mu(A)=\int_{A}fd\nu$ for all $A\in\mathcal{F}$. The function $f$ is the Radon-Nikodym derivative, denoted $\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$.
Metric. Let $\phi:\mathbb{R}_{\geq 0}\to\mathbb{R}$ be convex with $\phi(1)=0$. For probability measures $\mu\ll\nu$ on $(E,\mathcal{F})$, define the $\phi$-divergence $D_{\phi}(\mu||\nu)=\int_{E}\phi\!\left(\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\right)\mathrm{d}\nu$. For $\phi(x)=x\log x$ this is the Kullback-Leibler divergence, and for $\phi(x)=x^{2}-1$ it is the $\chi^{2}$-divergence. For $q>0$, the $q$-Rényi divergence is $\mathcal{R}_{q}(\mu||\nu)=\frac{1}{q-1}\log\!\bigl(\chi^{q}(\mu||\nu)+1\bigr)$. The relative Fisher information is $\mathrm{FI}(\mu||\nu)=\int_{E}\left\lVert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\right\rVert^{2}d\mu$.
Volumes, distance to a set and minwidth. Let $\mathsf{vol}(K)$ and $\mathsf{vol}_{d-1}(\partial K)$ respectively denote the volumes of $K\subseteq\mathbb{R}^{d}$ and the boundary set $\partial K\subseteq\mathbb{R}^{d-1}$. Let $d(y,K):=\inf_{z\in S}\|y-z\|$ denote the Euclidean distance from $y$ to $S$. Set $\mbox{minwidth}(K)=\min_{\left\lVert a\right\rVert=1}\left(\max_{y\in K}a^{T}y-\min_{y\in K}a^{T}y\right)$.
Normalizing constants. For a measurable $\Theta:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$, define the normalizing constant $N_{\Theta}:=\int_{\mathbb{R}^{m}}e^{-\Theta(z)}dz$ whenever this integral is finite.
Isoperimetric inequalities and isoperimetric constants. We say $\nu$ satisfies the log-Sobolev inequality (LSI) with constant $C_{\mathrm{LSI}}$ if for all $\mu\ll\nu$, $\mathrm{KL}(\mu||\nu)\leq\frac{C_{\mathrm{LSI}}}{2}\mathrm{FI}(\mu||\nu)$. We say $\nu$ satisfies the Poincaré inequality (PI) with constant $C_{\mathrm{PI}}$ if for any smooth bounded $\psi$, $\mathsf{Var}_{\nu}(\psi)\leq C_{\mathrm{PI}}\mathbb{E}_{\nu}\!\left[\left\lVert\nabla\psi\right\rVert^{2}\right]$.
In [41, Appendix C], the authors summarize recent bounds on isoperimetric constants [9, 35, 51, 12, 38, 39]. As a consequence, we obtain the following bounds on the LSI and PI constants for the uniform law on a convex body $K$. Recall that a probability measure $\pi$ on $K$ is isotropic if, for $X\sim\pi$, we have $\mathbb{E}X_{i}=0$ and $\mathbb{E}[X_{i}X_{j}]=\mathbf{1}_{i=j}$ for all $i,j$.
@@
[41, Lemma 18] Let $\pi$ be the uniform distribution over $K$ and $K\subset\mathbb{R}^{d}$ be a convex body with diameter $D$, where $D=\max_{x,y\in K}\left\lVert x-y\right\rVert$. Then we have $C_{\mathrm{PI}}(\pi)=\mathcal{O}\left(\left\lVert\mathrm{Cov}{\left(\pi\right)}\right\rVert_{\operatorname{op}}\log d\right)$ and $C_{\mathrm{LSI}}(\pi)=\mathcal{O}\left(D^{2}\right)$. In particular, if $\pi$ is isotropic, then $C_{\mathrm{PI}}(\pi)=\mathcal{O}\left(\log d\right)$ and $C_{\mathrm{LSI}}(\pi)=\mathcal{O}\left(D\right)$.
Oracles. A membership oracle decides whether a query $x\in\mathbb{R}^{d}$ lies in $K$. A separation oracle either certifies $x\in K$ or, if $x\notin K$, returns $g:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\langle g(x)\,,x-y\rangle\geq 0$ for all $y\in K$; in particular, it subsumes membership. A projection oracle returns $\operatorname{proj}_{K}(y)=\mathrm{argmin}\,\{\left\lVert x-y\right\rVert^{2}:x\in K\}$; clearly $\operatorname{proj}_{K}(y)=y$ if $y\in K$.
Warmness. For probability measures $\mu,\nu$ on $\mathbb{R}^{d}$ and $M>0$, we say $\mu$ is $M$-warm w.r.t. $\nu$ if $\mu\ll\nu$ and $\frac{\mathrm{d}\mu}{\mathrm{d}\nu}(x)\leq M$ for all $x\in\mathbb{R}^{d}$. In Algorithms 3 and 4, we assume a warm start: $\mu_{0}$ is $M$-warm w.r.t. the uniform distribution $\pi$ on $K$.
Standing assumptions for the paper. In both Sections 3 and 4, we assume the following conditions hold:
@@
    the initial distribution $\mu_{0}$ is $M$-warm with respect to the uniform distribution $\pi$ on $K$, i.e., $\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\pi}\leq M$.
2.2 The proximal sampler
The proximal sampler is first proposed by [48] to sample log-concave distribution in $\mathbb{R}^{d}$. Given a stepsize $\eta>0$, it aims to sample the target distribution $\pi^{X}(x)\sim\exp(-f(x))$ by performing Gibbs sampling for the augmented distribution $\pi^{X,Y}(x,y)\sim\exp\left(-f(x)-\left\lVert x-y\right\rVert^{2}/(2\eta)\right)$ whose $X$-marginal is the target $\pi^{X}$. This idea of sampling from a joint distribution to obtain the marginal distribution has been observed in earlier references, for example [16]. Each iteration of the proximal sampler alternates between two steps:
Algorithm 1 Proximal sampler by [48]